OpenStudy (anonymous):

Is there a point of oblique inflection at y=x^3 + 4x^2-9 Please show your working.

OpenStudy (anonymous):

\[\huge y=x^3 + 4x^2-9\] \[\huge \frac{dy}{dx}=3x^2+8x\] \[\huge \frac{d^2y}{dx^2}=6x+8\] At IP's (Inflection/Inflexion Points) \[\large \frac{d^2y}{dx^2}=0\] \[\huge 6x=-8\] \[\huge x=-\frac{4}{3}\] \[\huge y=-\frac{115}{27}\]